# A Comprehensive Study on Large Language Models for Mutation Testing

### v3

@@ 3.1. RQ1: Performance on Effectiveness
Coupling Rate: The overall Coupling Rate for each mutation generation approach is shown in the last row of Table 4. The traditional approaches, PIT, Major, LEAM, and Bert, exhibit lower coupling rates than LLM-based approaches, with coupling rates of 23.5%, 38.9%, 38.1%, and 47.3%, respectively. Among the LLM-based approaches, GPT-4o achieves the highest coupling rate, with a value of 50.6%. It is followed closely by GPT-4o-Mini and DeepSeek, with coupling rates of 49.5% and 49.5%, respectively. Overall, the trends are similar to the behavior similarity discussed in Section 3.1, indicating the superiority of the LLM-based approaches, particularly GPT-4o, in generating highly coupled mutations that better mimic real bugs.
| Category | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| Strong | 10.1% | 9.6% | 4.9% | 7.8% | 7.0% | 6.8% | 6.2% | 6.4% | 8.0% | 6.8% |
| Strong Addi. | 4.7% | 5.8% | 4.3% | 5.2% | 6.7% | 10.2% | 12.1% | 10.3% | 5.7% | 6.0% |
| Partial | 1.8% | 1.1% | 0.5% | 0.8% | 1.2% | 1.2% | 0.8% | 1.0% | 2.7% | 0.9% |
| Partial Addi. | 6.9% | 22.4% | 28.4% | 33.5% | 32.0% | 32.5% | 30.4% | 31.8% | 17.6% | 31.9% |
| Not Subs. | 30.6% | 13.2% | 26.7% | 20.4% | 23.3% | 23.1% | 21.3% | 21.9% | 16.7% | 21.3% |
| Not Detected | 45.9% | 47.9% | 35.2% | 32.3% | 29.8% | 26.3% | 29.2% | 28.6% | 49.3% | 33.1% |
@@
Coupling Categorization: As mentioned in Section 2.4.1, we classify each mutation of each generation approach into different scales, i.e., Strong Substitution, Strong + Additional Tests Fail, Partial Substitution, Partial + Additional Tests Fail, Not Substitution, and Not Detected. Note that the first four categories are used to calculate the Coupling Rate. Table 7 presents the distribution of each generation approach. Among the traditional approaches, PIT and Major exhibit a significantly higher proportion of mutants in the Not Detected category, with rates of 45.9% and 47.9%, respectively, indicating that a substantial number of their generated mutants fail to couple with real bugs. However, they achieve relatively higher proportions in the Strong category, with PIT at 10.1% and Major at 9.6%, suggesting that while their coupling ability is limited, they can still generate a small number of strongly coupled mutations. In contrast, LLM-based approaches such as GPT-4o and GPT-4o-Mini, demonstrate stronger coupling capabilities across multiple categories. These approaches achieve significantly higher rates in the Partial Substitution + Additional Tests Fail category, with values of 32.5% and 30.4%, respectively. Furthermore, they outperform traditional methods in the Strong Substitution + Additional Tests Fail category, with GPT-4o and GPT-4o-Mini achieving 10.2% and 12.1%, respectively. Overall, while the traditional rule-based approaches generate a larger proportion of strong substitution mutations, LLM-based approaches demonstrate a stronger ability to generate diverse and highly coupled mutations.
@@
AST Node Diversity: To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element). For the non-deletion mutations, we parse the code before and after mutation by the parser Javalang333https://pypi.org/project/javalang, to observe which new AST node types are introduced. A larger number of newly introduced AST node types indicates richer diversity.
@@
Figure 2 illustrates the number of introduced AST node types, with traditional approaches highlighted in green and LLM-based approaches in blue. The figure reveals that LLM-based mutation approaches introduce more types of new AST nodes than traditional approaches. The best LLM-based approach, DeepSeek, exhibits the greatest diversity, introducing 49 new AST node types, followed by GPT-4o-Mini, which introduces 45 new types. Additionally, all six LLMs are significantly higher than all other approaches in AST node types. We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 3), as shown by the “Examples” in Figure 2. Interestingly, we find that these examples introduce only one new type of AST node and are subsumed by all other approaches. This indicates that LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
| Top K | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| Del. | 15.3% | 48.9% | 0.4% | 5.3% | 2.1% | 0.5% | 1.0% | 0.3% | 0.1% |
| 1st | ST(78.6%) | LT(23.2%) | LT(29.4%) | LT(39.2%) | LT(45.4%) | LT(43.3%) | LT(38.8%) | LT(22.7%) | LT(47.8%) |
| 2nd | LT(21.4%) | MR(14.4%) | SE(25.7%) | BO(17.3%) | BO(16.6%) | BO(21.6%) | BO(19.1%) | BO(13.8%) | BO(25.7%) |
| 3rd | —— | CT(9.6%) | MR(24.5%) | MI(9.9%) | CC(8.8%) | CC(7.6%) | MI(10.1%) | MR(13.1%) | MI(9.3%) |
@@
Del.: The proportion of deletion mutations. LT: Literal, BO: BinaryOperation, MI: MethodInvocation, MR: MemberReference, SE: StatementExpression, ST: Statement.
@@
We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 8, which highlights the top three most frequent AST node types. For the proportion of deletions, the traditional approach, LEAM, shows the highest deletion mutation ratio, with 48.9%, followed by Major at 15.3%. In contrast, all LLM-based approaches generate deletion mutations at a rate below 6%, with three models producing less than 1.0%. This indicates that, compared to existing approaches, LLMs are less inclined to generate deletion mutations. As shown later, we find that deletions are one of the key factors contributing to the lower compilation success rate.
@@
| Distance | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| 1 | 69.2% | 58.5% | 73.8% | 52.4% | 50.0% | 47.3% | 45.1% | 36.0% | 46.4% |
| 2 | 27.8% | 23.6% | 8.9% | 36.4% | 38.8% | 41.9% | 44.2% | 39.1% | 46.6% |
| 3 | 1.8% | 7.3% | 2.7% | 6.8% | 7.2% | 6.3% | 6.5% | 8.0% | 3.8% |
| 4 | 0.6% | 3.4% | 14.2% | 2.4% | 2.2% | 2.5% | 2.6% | 6.0% | 1.8% |
| 5 | 0.4% | 1.0% | 0.1% | 0.9% | 0.8% | 0.8% | 0.8% | 3.7% | 0.5% |
| 6 | 0.2% | 6.1% | 0.4% | 1.0% | 1.0% | 1.2% | 0.8% | 7.2% | 1.0% |
From the table, we can find that all traditional approaches predominantly generate small-scale changes. Specifically, the proportion of mutations with an edit distance of 1 for Major, LEAM, and Bert is 69.2%, 58.5%, and 73.8%, respectively, indicating they involve minimal syntactic changes to the original code in the majority of cases. In contrast, LLM-based approaches exhibit a broader range of edit distances, reflecting their capability to introduce more complex transformations. For example, GPT-4o exhibits a more balanced distribution, with 38.8% of its mutations having an edit distance of 2, 7.2% with a distance of 3, and 2.2% with a distance of 4, indicating it generates both minor and moderate code changes.
@@ 3.2. RQ2: Performance on Validity
To evaluate the validity of mutation generation approaches, we consider the aspects that may impact their validity, i.e., comparability rate, duplicate mutation rate, and equivalent mutation rate. This RQ is divided into the following sub-RQs.
@@
The row Comp. Rate of Table 4 presents the overall Compilability Rate (CR) of each approach. Among the traditional approaches, the rule-based approaches, PIT and Major, demonstrate superior compilability, with PIT achieving a perfect 100% , and Major following closely at 97.60%. Note that although Major operates under simple syntax rules, it can still generate non-compilable mutants. For example, the Java compiler’s analyzer rejects constructs like while(true). PIT uses the mutation operators more restrictedly and previously filters out non-compilable mutations. In contrast, LEAM and Bert, have rates at 35.0% and 22.5%, respectively. For LLM-based approaches, there is a considerable variation in compilability. GPT-4o achieves the highest compilability among LLMs at 76.40%, followed closely by GPT-4o-Mini and DeepSeek, with rates of 75.30% and 75.80%, respectively. Other LLMs such as GPT-3.5 and CodeLlama show moderate s of 62.50% and 71.20%, respectively, whereas models like StarChat fall behind with a of 38.50%. Compared to small models (i.e., LEAM and Bert), GPTs and DeepSeek show significantly higher compilation success rates. For example, the compilability rate of GPT-4o is over 30% higher than that of LEAM and Bert.
@@
As aforementioned, duplicate mutations refer to those that are redundant or identical to the original code, providing no meaningful contribution to the mutation testing process. Therefore, a lower Duplicate Mutation Rate () is desirable, as it indicates a more efficient mutation generation process with minimal redundancy. The row Dup. Mut. Rate of Table 4 presents the results.
Rule-based approaches, PIT and Major, achieve a perfect 0% . This is expected, as these approaches apply well-defined mutation rules that ensure each generated mutant introduces a meaningful change to the program. Other traditional approaches, i.e., LEAM and Bert, show relatively low s of 2.30% and 1.70%, respectively. In contrast, LLMs generate a significant portion of duplicate mutations, exhibiting a higher . For the LLMs from OpenAI, GPT-3.5, GPT-4o, and GPT-4o-Mini exhibit moderate s of 10.30%, 7.80%, and 7.00%, respectively. The open-source LLMs such as StarCoder and DeepSeek demonstrate higher s of 12.80% and 7.70%, respectively, while CodeLlama shows the highest at 37.10%.
@@
In total, 959 mutations were labeled, 508 of which were filtered by tests (i.e., killed), and the remaining 451 were manually checked. The entire process took approximately ten hours per person. The manual verification requires checking the code context, which may involve multiple classes and methods. For example, to check a mutation that replaces an integer with a method invocation, where the method is invoked by the this object but defined in the parent class, we should review the parent class file and verify whether its return value is consistently equal to the integer. We measure the consistency between the two authors by Cohen’s kappa coefficient () (Viera et al., 2005), achieving an overall value of 0.8, which indicates a high level of agreement. In cases of disagreement, both authors conducted thorough discussions to resolve contradictions and finally reach consistent labels. The labeling results are available at our repo.
| Metric | PIT | Major | LEAM | Bert | GPT-3.5 | GPT-4o | GPT-4o-Mini | DS-236b | SC-16b | CL-13b |
| #Sampled | 97 | 98 | 96 | 94 | 96 | 96 | 96 | 96 | 94 | 96 |
| #Test Filtered | 26 | 52 | 48 | 53 | 60 | 64 | 59 | 63 | 32 | 51 |
| #Eq. Mut | 1 | 2 | 2 | 6 | 4 | 3 | 3 | 5 | 10 | 6 |
| EMR | 1.0% | 2.0% | 2.1% | 6.4% | 4.2% | 3.1% | 3.1% | 5.2% | 10.6% | 6.3% |
Table 10 presents the estimated Equivalent Mutation Rate (), reporting the number of sampled mutations (#Sampled), the number of equivalent mutations we identified (#Eq. Mut), and the for each approach. Among the traditional approaches, PIT achieves the lowest of 1.0%, followed by Major with an of 2.0%, indicating that rule-based techniques produce fewer equivalent mutations. In contrast, the LLM-based approaches exhibit higher s, with variations across different models. StarChat has the highest s of 10.6%, while GPT-4o, GPT-4o-Mini, and DeepSeek show relatively lower s, ranging from 3.1% to 5.2%. Overall, although the OpenAI models such as GPT-4o and GPT-4o-Mini, tend to have slightly higher s than PIT and Major, the increase remains modest and within an acceptable range.
Note that the estimated of Major given by the recent study of Kushigian et al. (Kushigian et al., 2024) is 2.97%, where the authors sampled a subset of mutation for manual judgment. We adopt the same experimental process and the result of Major is 2.0%, which is consistent with theirs considering the randomness introduced by sampling.

### v4

@@ 3.1. RQ1: Performance on Effectiveness
BF: BugFarm. LLMph: LLMorpheus. LLMut: our prompt. Moc.: Mockito. Clo.: Closure. Co.: Codec. Gs.: Gson. Ja.: JacksonCore. Js.: Jsoup. ConD: ConDefects.
Coupling Rate: The overall Coupling Rate for each mutation generation approach is shown in the last row of Table 4. The traditional approaches, PIT, Major, LEAM, and BERT, exhibit lower coupling rates than LLM-based approaches, with coupling rates of 23.5%, 38.9%, 38.1%, and 47.3%, respectively. Among the LLM-based approaches, LLMorpheus achieves the highest coupling rate, with a value of 51.7%. It is followed closely by DeepSeek-V3-761b and GPT-4o-Mini based on our prompt, with coupling rates of 50.8% and 50.6%, respectively. Overall, the trends are similar to the behavior similarity discussed in Section 3.1, indicating the superiority of the LLM-based approaches, particularly LLMorpheus, DeepSeek-V3-761b, and GPT-4o, in generating overall highly coupled mutations that better mimic real bugs.
| Category | Strong | Strong Addi | Partial | Partial Addi | Not Subs. | Not Detected |
| PIT | 10.1% | 4.7% | 1.8% | 6.9% | 30.6% | 45.9% |
| Major | 9.6% | 5.8% | 1.1% | 22.4% | 13.2% | 47.9% |
| LEAM | 4.9% | 4.3% | 0.5% | 28.4% | 26.7% | 35.2% |
| BERT | 7.8% | 5.2% | 0.8% | 33.5% | 20.4% | 32.3% |
| BugFarm (GPT-4o) | 8.4% | 5.4% | 0.6% | 33.6% | 28.3% | 23.8% |
| LLMorpheus (GPT-4o) | 9.3% | 4.9% | 2.1% | 35.4% | 23.6% | 24.6% |
| LLMut (GPT-3.5) | 7.0% | 6.7% | 1.2% | 32.0% | 23.3% | 29.8% |
| LLMut (GPT-4o) | 6.8% | 10.2% | 1.2% | 32.5% | 23.1% | 26.3% |
| LLMut (GPT-4o-M) | 6.2% | 12.1% | 0.8% | 30.4% | 21.3% | 29.2% |
| LLMut (DS-761b) | 11.2% | 4.6% | 1.9% | 33.1% | 24.1% | 25.1% |
| LLMut (DS-236b) | 6.4% | 10.3% | 1.0% | 31.8% | 21.9% | 28.6% |
| LLMut (SC-16b) | 8.0% | 5.7% | 2.7% | 17.6% | 16.7% | 49.3% |
| LLMut (CL-13b) | 6.8% | 6.0% | 0.9% | 31.9% | 21.3% | 33.1% |
@@
Coupling Categorization: As mentioned in Section 2.4.1, we classify each mutation of each generation approach into different scales, i.e., Strong Substitution, Strong + Additional Tests Fail, Partial Substitution, Partial + Additional Tests Fail, Not Substitution, and Not Detected. Note that the first four categories are used to calculate the Coupling Rate, indicating these mutations semantically overlap with their corresponding real bugs. Table 7 presents the distribution of each generation approach. Among the traditional approaches, PIT and Major exhibit a significantly higher proportion of mutants in the Not Detected category, with rates of 45.9% and 47.9%, respectively, indicating that a substantial number of their generated mutants fail to couple with real bugs. However, they achieve relatively higher proportions in the Strong category, with PIT at 10.1% and Major at 9.6%, suggesting that while their coupling ability is limited, they can still generate a small number of strongly coupled mutations. In contrast, LLM-based approaches such as DeepSeek-V3-761b, GPT-4o, and GPT-4o-Mini, demonstrate stronger coupling capabilities across multiple categories. For instance, DeepSeek-V3-761b achieves the highest proportion in the Strong category at 11.2%, while GPT-4o-Mini leads in the Strong + Additional Tests Fail category, with 12.1%. Overall, while the traditional rule-based approaches generate a larger proportion of strong substitution mutations, LLM-based approaches demonstrate a stronger ability to generate diverse and highly coupled mutations.
@@
AST Node Diversity: To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element). For the non-deletion mutations, we parse the code before and after mutation by the parser Javalang555https://pypi.org/project/javalang, to observe which new AST node types are introduced. A larger number of newly introduced AST node types indicates richer diversity.
@@
Figure 2 illustrates the number of introduced AST node types, with traditional approaches highlighted in green and LLM-based approaches in blue. The figure reveals that LLM-based mutation approaches introduce more types of new AST nodes than traditional approaches. The best LLM-based approaches, DeepSeek-V3-761b and DeepSeek-Coder-V2-236b, exhibit the greatest diversity, introducing 49 new AST node types, followed by GPT-4o-Mini, which introduces 45 new types. Additionally, all six LLMs based on our prompt are significantly higher than all other approaches in AST node types. The existing LLM-based approaches, BugFarm and LLMorpheus, introduce 31 and 21 new AST node types, respectively. While BugFarm outperforms traditional approaches, it still lags behind ours. We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 2), as shown by the “Examples” in Figure 2. Interestingly, we find that these examples introduce only one new type of AST node and are subsumed by all other approaches. This indicates that LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
| Approach | Deletion | 1st | 2nd | 3rd |
| Major | 15.3% | ST (78.6%) | LT (21.4%) | —— |
| LEAM | 48.9% | LT (23.2%) | MR (14.4%) | CT (9.6%) |
| BERT | 0.4% | LT (29.4%) | SE (25.7%) | MR (24.5%) |
| BugFarm (GPT-4o) | 5.5% | LT (41.9%) | BO (12.2%) | MI (9.0%) |
| LLMorpheus (GPT-4o) | 0.3% | LT (24.0%) | BO (23.7%) | MR (15.8%) |
| LLMut (GPT-3.5) | 5.3% | LT (39.2%) | BO (17.3%) | MI (9.9%) |
| LLMut (GPT-4o) | 2.1% | LT (45.4%) | BO (16.6%) | CC (8.8%) |
| LLMut (GPT-4o-M) | 0.5% | LT (43.3%) | BO (21.6%) | CC (7.6%) |
| LLMut (DS-761b) | 3.8% | LT (44.2%) | BO (15.6%) | CC (8.6%) |
| LLMut (DS-236b) | 1.0% | LT (38.8%) | BO (19.1%) | MI (10.1%) |
| LLMut (SC-16b) | 0.3% | LT (22.7%) | BO (13.8%) | MR (13.1%) |
| LLMut (CL-13b) | 0.1% | LT (47.8%) | BO (25.7%) | MI (9.3%) |
@@
Deletion: The proportion of deletion mutations. BO: BinaryOperation. CC: ClassCreator. CT: Cast. LT: Literal. MI: MethodInvocation. MR: MemberReference. SE: StatementExpression. ST: Statement.
@@
We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 8, which highlights the proportion of the mutations deleting the target code element and the top three most frequent AST node types. For the proportion of deletions, the traditional approach, LEAM, shows the highest deletion mutation ratio, with 48.9%, followed by Major at 15.3%. In contrast, all LLM-based approaches generate deletion mutations at a rate below 6%, with three models producing less than 1.0%. This indicates that, compared to existing approaches, LLMs are less inclined to generate deletion mutations. As shown later, we find that deletions are one of the key factors contributing to the lower compilation success rate.
@@
| Distance | 1 | 2 | 3 | 4 | 5 | 6 |
| Major | 69.2% | 27.8% | 1.8% | 0.6% | 0.4% | 0.2% |
| LEAM | 58.5% | 23.6% | 7.3% | 3.4% | 1.0% | 6.1% |
| BERT | 73.8% | 8.9% | 2.7% | 14.2% | 0.1% | 0.4% |
| BugFarm (GPT-4o) | 44.7% | 43.5% | 8.3% | 1.6% | 0.6% | 1.4% |
| LLMorpheus (GPT-4o) | 54.2% | 25.7% | 14.6% | 3.7% | 0.9% | 0.8% |
| LLMut (GPT-3.5) | 52.4% | 36.4% | 6.8% | 2.4% | 0.9% | 1.0% |
| LLMut (GPT-4o) | 50.0% | 38.8% | 7.2% | 2.2% | 0.8% | 1.0% |
| LLMut (GPT-4o-M) | 47.3% | 41.9% | 6.3% | 2.5% | 0.8% | 1.2% |
| LLMut (DS-761b) | 46.1% | 44.2% | 6.8% | 1.5% | 0.4% | 0.9% |
| LLMut (DS-236b) | 45.1% | 44.2% | 6.5% | 2.6% | 0.8% | 0.8% |
| LLMut (SC-16b) | 36.0% | 39.1% | 8.0% | 6.0% | 3.7% | 7.2% |
| LLMut (CL-13b) | 46.4% | 46.6% | 3.8% | 1.8% | 0.5% | 1.0% |
From the table, we can find that all traditional approaches predominantly generate small-scale changes. Specifically, the proportion of mutations with an edit distance of 1 for Major, LEAM, and BERT is 69.2%, 58.5%, and 73.8%, respectively, indicating they involve minimal syntactic changes to the original code in the majority of cases. In contrast, LLM-based approaches exhibit a broader range of edit distances, reflecting their capability to introduce more complex transformations. For example, DeepSeek-V3-761b exhibits a more balanced distribution, with 44.2% of its mutations having an edit distance of 2, 7.2% with a distance of 3, and 2.2% with a distance of 4, indicating it generates both minor and moderate code changes.
@@ 3.2. RQ2: Performance on Validity
To evaluate the validity of mutation generation approaches, we consider the aspects that may impact their validity, i.e., compilability rate, duplicate mutation rate, and equivalent mutation rate. This RQ is divided into the following sub-RQs.
@@
The column Comp. Rate of Table 4 presents the overall Compilability Rate () of each approach. Among the traditional approaches, the rule-based approaches, PIT and Major, demonstrate superior compilability, with PIT achieving a perfect 100% , and Major following closely at 97.60%. Note that although Major operates under simple syntax rules, it can still generate non-compilable mutants. For example, the Java compiler’s analyzer rejects constructs like while(true). PIT uses the mutation operators more restrictively and previously filters out non-compilable mutations. In contrast, LEAM and BERT, have rates at 35.0% and 22.5%, respectively. For LLM-based approaches, there is a considerable variation in compilability. The existing LLM-based approaches, BugFarm and LLMorpheus, achieve relatively high rates of 74.1% and 67.8%, respectively. Among the LLMs using our prompt, DeepSeek-V3-761b attains the highest at 77.6%, followed closely by GPT-4o (76.4%), DeepSeek-Coder-V2-236b (75.8%), and GPT-4o-Mini (75.3%). Compared to small models (i.e., LEAM and BERT), GPTs and DeepSeek models show significantly higher compilation success rates. For example, the compilability rate of GPT-4o is over 30% higher than that of LEAM and BERT.
@@
As aforementioned, duplicate mutations refer to those that are redundant or identical to the original code, providing no meaningful contribution to the mutation testing process. Therefore, a lower Duplicate Mutation Rate () is desirable, as it indicates a more efficient mutation generation process with minimal redundancy. The column Dup. Mut. Rate of Table 4 presents the results.
Rule-based approaches, PIT and Major, achieve a perfect 0% . This is expected, as these approaches apply well-defined mutation rules that ensure each generated mutant introduces a meaningful change to the program. Other traditional approaches, i.e., LEAM and BERT, show relatively low s of 2.3% and 1.7%, respectively. In contrast, LLMs generate a significant portion of duplicate mutations, exhibiting a higher . For the LLMs from OpenAI, GPT-3.5, GPT-4o, and GPT-4o-Mini exhibit moderate s of 10.3%, 7.8%, and 7.0%, respectively. For the existing LLM-based approaches, BugFarm and LLMorpheus, exhibit similar s of 7.5% and 7.1%, respectively. The open-source LLMs such as StarCoder and DeepSeek-Coder-V2-236b demonstrate higher s of 12.8% and 7.7%, respectively, while CodeLlama shows the highest at 37.1%.
@@
In total, 1247 mutations were labeled, 709 of which were filtered by tests (i.e., killed), and the remaining 538 were manually checked. The entire process took approximately twelve hours per person. The manual verification requires checking the code context, which may involve multiple classes and methods. For example, to check a mutation that replaces an integer with a method invocation, where the method is invoked by the this object but defined in the parent class, we should review the parent class file and verify whether its return value is consistently equal to the integer. We measure the consistency between the two authors by Cohen’s kappa coefficient () (Viera et al., 2005), achieving an overall value of 0.8, which indicates a high level of agreement. In cases of disagreement, both authors conducted thorough discussions to resolve contradictions and finally reach consistent labels. The labeling results are available at our repo.
| Approach | #Sampled | #Test Filtered | #Eq. Mut. | EMR |
| PIT | 97 | 26 | 1 | 1.0% |
| Major | 96 | 52 | 2 | 2.1% |
| LEAM | 96 | 48 | 2 | 2.1% |
| BERT | 94 | 64 | 3 | 6.4% |
| BugFarm (GPT-4o) | 96 | 67 | 3 | 3.1% |
| LLMorpheus (GPT-4o) | 96 | 69 | 4 | 4.2% |
| LLMut (GPT-3.5) | 96 | 60 | 4 | 4.2% |
| LLMut (GPT-4o) | 96 | 64 | 3 | 3.1% |
| LLMut (GPT-4o-M) | 96 | 59 | 3 | 3.1% |
| LLMut (DS-761b) | 96 | 65 | 6 | 6.3% |
| LLMut (DS-236b) | 96 | 63 | 5 | 5.2% |
| LLMut (SC-16b) | 94 | 32 | 10 | 10.6% |
| LLMut (CL-13b) | 96 | 51 | 6 | 6.3% |
Table 10 presents the estimated Equivalent Mutation Rate (), reporting the number of sampled mutations (#Sampled), the number of equivalent mutations we identified (#Eq. Mut), and the for each approach. Among the traditional approaches, PIT achieves the lowest of 1.0%, followed by Major with an of 2.1%, indicating that rule-based techniques produce fewer equivalent mutations. In contrast, the LLM-based approaches exhibit higher s, with variations across different models. StarChat has the highest s of 10.6%, while GPT-4o, GPT-4o-Mini, and DeepSeek-Coder-V2-236b show relatively lower s, ranging from 3.1% to 5.2%. Overall, although the OpenAI models such as GPT-4o and GPT-4o-Mini tend to have slightly higher s than PIT and Major, the increase remains modest and within an acceptable range.
Note that the estimated of Major given by the recent study of Kushigian et al. (Kushigian et al., 2024) is 2.97%, where the authors sampled a subset of mutations for manual judgment. We adopt the same experimental process and the result of Major is 2.0%, which is consistent with theirs, considering the randomness introduced by sampling. Moreover, the definition of equivalent mutations in this paper is the set of syntactically different but semantically equivalent mutations, which may differ from existing studies (Tip et al., 2024).